find the equation of a line passing through the point 3,2 and perpendicular to the line x=y

Grade:11

1 Answers

Avani Patidar
8 years ago
the line is y=5-x
since the line is perpendicular to line y=x, the slope of this line is -1.
and acc. to point slope formula i.e., (y-y’)=m(x-x’)
hence (y-2)=(-1)(x-3)
y-2=3-x
y=5-x